# Translating predictive distributions into informative priors

### v3

@@ 4 Conclusion
One limitation of the current work is that we only partly address non-uniqueness, but we emphasise that our methodology remains valuable in such settings. Specifically, our approach provides insight into consequences of certain $\text{T}(Y\mid\boldsymbol{X})$: it facilitates identifying which components of $\lambda$ are well identified, and consideration of whether any difference between $\text{T}(Y\mid X_{r})$ and $\text{P}(Y\mid\lambda^{*},X_{r})$ are attributable to model inflexibility or an implausible target. We also have the opportunity to re-assess whether we have information that we could employ to fix certain components within $\lambda$ (e.g. the fixed prior for the noise in the human height example). Another limitation of our current work is that global optimisation methods lack guarantees of finding the global optimum in finite time, so the generalisability of the performance of our optimisation process requires further investigation in the future.
@@
We thank Daniela De Angelis and Mevin Hooten for their feedback; and the The Alan Turing Institute under the UK Engineering and Physical Sciences Research Council (EPSRC) [EP/N510129/1] and the UK Medical Research Council [MC_UU_00002/2 and MC_UU_00002/20] for support. No original data were generated; the fda package for R contains the growth data. R code for the examples is at https://gitlab.com/andrew-manderson/pbbo-paper. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ 9         Sample importance points $(y_{i,r})_{i=1}^{I_{r}}\sim\text{Q}(Y\mid X_{r})$
$\mathcal{Y}$ $\text{Q}_{r}(Y)$ Parameter estimates Mixture weights
Notes
$\mathbb{R}$ $&\pi_{1}\text{Student-}t_{5}(Y;\hat{\mu}_{1},\hat{s}_{1})+\\ &\pi_{2}\text{Student-}t_{5}(Y;\hat{\mu}_{2},\hat{s}_{2})\end{aligned}$ $\hat{\mu}_{1}=\hat{\mu}^{(\text{P})},&\,\hat{s}_{1}=c\sqrt{\hat{v}^{(\text{P})}}\\ \hat{\mu}_{2}=\hat{\mu}^{(\text{T})},&\,\hat{s}_{2}=c\sqrt{\hat{v}^{(\text{T})}}\\ \end{aligned}$ $\pi_{1}=\pi_{2}=0.5$
@@
$(0,\infty)$ $&\pi_{1}\text{Gamma}(Y;\hat{\alpha}_{1},\hat{\beta}_{1})+\\ &\pi_{2}\text{Gamma}(Y;\hat{\alpha}_{2},\hat{\beta}_{2})\end{aligned}$ $\hat{\alpha}_{1}=\frac{(\hat{\mu}^{(\text{P})})^{2}}{\tilde{\omega}^{(\text{P})}},&\,\hat{\beta}_{1}=\frac{\hat{\mu}^{(\text{P})}}{\tilde{\omega}^{(\text{P})}}\\ \hat{\alpha}_{2}=\frac{(\hat{\mu}^{(\text{T})})^{2}}{\tilde{\omega}^{(\text{T})}},&\,\hat{\beta}_{2}=\frac{\hat{\mu}^{(\text{T})}}{\tilde{\omega}^{(\text{T})}}\\ \end{aligned}$ $\pi_{1}=\pi_{2}=0.5$
$\tilde{\omega}=\min(c^{2}\hat{v},10^{5})$,
$c$ defaults to $1.05$
$[0,a]$ $&\frac{\pi_{1}}{a}\text{Beta}\left(\frac{Y}{a};\hat{a}_{1},\hat{b}_{1}\right)+\\ &\frac{\pi_{2}}{a}\text{Beta}\left(\frac{Y}{a};\hat{a}_{2},\hat{b}_{2}\right)+\\ &\pi_{3}\mathbbm{1}_{\{Y=a\}}\end{aligned}$ $\hat{a}_{1}&=\hat{\mu}^{(\text{P})}\left[\frac{\hat{\mu}^{(\text{P})}}{\tilde{\omega}^{(\text{P})}}(1-\hat{\mu}^{(\text{P})})-1\right]\\ \hat{b}_{1}&=\frac{(1-\hat{\mu}^{(\text{P})})}{\hat{\mu}^{(\text{P})}}\hat{a}_{1}\\ \hat{a}_{2}&=\hat{\mu}^{(\text{T})}\left[\frac{\hat{\mu}^{(\text{T})}}{\tilde{\omega}^{(\text{T})}}(1-\hat{\mu}^{(\text{T})})-1\right]\\ \hat{b}_{2}&=\frac{(1-\hat{\mu}^{(\text{T})})}{\hat{\mu}^{(\text{T})}}\hat{a}_{2}\end{aligned}$ $\begin{gathered}\pi_{1}=\pi_{2}=0.45\\ \pi_{3}=0.05\end{gathered}$
$\tilde{\omega}=\max(c^{2}\hat{v},10^{-6})$,
@@
Lastly, Algorithm 5 summaries the entire methodology we introduce in this paper.
@@ 7         Update design $\mathcal{D}$ using $\mathcal{P}_{b}$ and $\mathcal{D}_{b}$ via Algorithm 3
S4 Additional information for the cure fraction survival example
@@
S4.1 Hyperparameter support $\Lambda$
@@
S4.2 Pareto Frontiers for Cramér-von Mises discrepancy
@@
S4.3 Final objective values
@@
S5 Additional information for the $R^{2}$ example
S5.1 Hyperparameter support $\Lambda$ – faithfulness experiment
@@
S5.2 Full faithfulness results
@@
Figure 17: As in Figure 7 but for all values of $(s_{1},s_{2})$ denoted in the facet panels titles. The performance of the regularised horseshoe is superior to the Dirichlet-Laplace, both of which are vast improvements over the Gaussian.
S5.3 A comparison to an asymptotic result
@@
S5.3.1 Results and analytic comparison
@@
S6 Additional information for the Preece-Baines example
S6.1 Hyperparameter support $\Lambda$
@@
S6.2 Details for $\text{T}(Y)$ and $\text{T}(Y\mid X_{r})$
@@
S6.4 Pareto frontiers for the covariate-independent target
The Pareto frontier for the covariate-independent target and all values of $\kappa\in\mathcal{K}$ is displayed in Figure 21.
@@
S6.5 Final predictive discrepancy values for the human growth example
The located optimal values of predictive discrepancy are displayed in Figure 22.
@@
Figure 22: Final predictive discrepancy $\log(D(\lambda^{*}))$, or $\log(D(\lambda^{*}\mid\boldsymbol{X}))$ for the covariate-specific target. The multiple objective optimisation approach uses the previously selected values of $\kappa=0.15$ for the covariate-independent target, and $\kappa=0.2$ for the covariate-specific target. Horizontal jitter has been applied for readability.
S6.6 Full marginal prior and posterior comparison plots
@@
Figures 23 and 24 are extended versions of Figure 14 in the main text, and display the prior and posterior estimates for all the parameters in $\theta$. Consistency and uniqueness remain, evidently, challenging and as yet unobtainable.
@@
Figure 23: A comparison of the priors (blue) produced by our method using the covariate-independent marginal target (bottom two rows); and Hartmann et al. (2020) (second row), with no prior displayed for the flat prior scenario. The corresponding posteriors (red) for individual $n=26$ under each of these priors are displayed as dashed lines. Note that y-axes change within columns and are limited to values that clip some of the priors/posteriors for readability.

### v4

@@ 4 Conclusion
One limitation of the current work is that we only partly address non-uniqueness, but we emphasise that our methodology remains valuable in such settings. Specifically, our approach provides insight into consequences of certain $\text{T}(Y\mid\boldsymbol{X})$: it facilitates discovering which components of $\lambda$ are uniquely determined, and consideration of whether any differences between $\text{T}(Y\mid X_{r})$ and $\text{P}(Y\mid\lambda^{*},X_{r})$ are attributable to model inflexibility or an implausible target. We also have the opportunity to re-assess whether we have information that we could employ to fix certain components within $\lambda$ (e.g. the fixed prior for the noise in the human height example). Another limitation of our current work is that global optimisation methods lack guarantees of finding the global optimum in finite time: results for CRS2 are largely empirical and results for multi-objective Bayesian optimisation remain a topic of research (e.g. Chowdhury & Gopalan, 2021). The generalisability of our optimisation process thus requires further investigation. Finally, the choice of secondary objective also invites future investigation into alternatives: practitioners may have other principles they wish to encode into the prior-setting process. Alternative objectives that instead minimise the variation in only a subset of $\theta$ whilst maximising the remaining parameters, or objectives that are functions of the joint distribution of $\theta$, are avenues for further research.
@@
We thank Daniela De Angelis and Mevin Hooten for their feedback; and the The Alan Turing Institute under the UK Engineering and Physical Sciences Research Council (EPSRC) [EP/N510129/1] and the UK Medical Research Council [MC_UU_00002/2, MC_UU_00002/20 and MC_UU_00040/04] for support. No original data were generated; the fda package for R contains the growth data. R code for the examples is at https://gitlab.com/andrew-manderson/pbbo-paper. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ 9         Sample importance points $(y_{i,r})_{i=1}^{I_{r}}\sim\text{Q}(Y\mid X_{r})$
$\mathcal{Y}$ $\text{Q}_{r}(Y)$ Parameter estimates Mixture weights Notes
$\mathbb{R}$ $&\pi_{1}\text{Student-}t_{5}(Y;\hat{\mu}_{1},\hat{s}_{1})+\\ &\pi_{2}\text{Student-}t_{5}(Y;\hat{\mu}_{2},\hat{s}_{2})\end{aligned}$ $\hat{\mu}_{1}=\hat{\mu}^{(\text{P})},&\,\hat{s}_{1}=c\sqrt{\hat{v}^{(\text{P})}}\\ \hat{\mu}_{2}=\hat{\mu}^{(\text{T})},&\,\hat{s}_{2}=c\sqrt{\hat{v}^{(\text{T})}}\\ \end{aligned}$ $\pi_{1}=\pi_{2}=0.5$ $c$ defaults to $1.05$
$(0,\infty)$ $&\pi_{1}\text{Gamma}(Y;\hat{\alpha}_{1},\hat{\beta}_{1})+\\ &\pi_{2}\text{Gamma}(Y;\hat{\alpha}_{2},\hat{\beta}_{2})\end{aligned}$ $\hat{\alpha}_{1}=\frac{(\hat{\mu}^{(\text{P})})^{2}}{\tilde{\omega}^{(\text{P})}},&\,\hat{\beta}_{1}=\frac{\hat{\mu}^{(\text{P})}}{\tilde{\omega}^{(\text{P})}}\\ \hat{\alpha}_{2}=\frac{(\hat{\mu}^{(\text{T})})^{2}}{\tilde{\omega}^{(\text{T})}},&\,\hat{\beta}_{2}=\frac{\hat{\mu}^{(\text{T})}}{\tilde{\omega}^{(\text{T})}}\\ \end{aligned}$ $\pi_{1}=\pi_{2}=0.5$ $\tilde{\omega}=\min(c^{2}\hat{v},10^{5})$,
@@
$[0,a]$ $&\frac{\pi_{1}}{a}\text{Beta}\left(\frac{Y}{a};\hat{a}_{1},\hat{b}_{1}\right)+\\ &\frac{\pi_{2}}{a}\text{Beta}\left(\frac{Y}{a};\hat{a}_{2},\hat{b}_{2}\right)+\\ &\pi_{3}\mathbbm{1}_{\{Y=a\}}\end{aligned}$ $\hat{a}_{1}&=\hat{\mu}^{(\text{P})}\left[\frac{\hat{\mu}^{(\text{P})}}{\tilde{\omega}^{(\text{P})}}(1-\hat{\mu}^{(\text{P})})-1\right]\\ \hat{b}_{1}&=\frac{(1-\hat{\mu}^{(\text{P})})}{\hat{\mu}^{(\text{P})}}\hat{a}_{1}\\ \hat{a}_{2}&=\hat{\mu}^{(\text{T})}\left[\frac{\hat{\mu}^{(\text{T})}}{\tilde{\omega}^{(\text{T})}}(1-\hat{\mu}^{(\text{T})})-1\right]\\ \hat{b}_{2}&=\frac{(1-\hat{\mu}^{(\text{T})})}{\hat{\mu}^{(\text{T})}}\hat{a}_{2}\end{aligned}$ $\begin{gathered}\pi_{1}=\pi_{2}=0.45\\ \pi_{3}=0.05\end{gathered}$ $\tilde{\omega}=\max(c^{2}\hat{v},10^{-6})$,
@@
Lastly, Algorithm 5 summarises the entire methodology we introduce in this paper.
@@ 7         Update design $\mathcal{D}$ using $\mathcal{P}_{b}$ and $\mathcal{D}_{b}$ via Algorithm 3
S4 Using the Kullback–Leibler divergence as a discrepancy
Our choice of discrepancy is general but arbitrary. Another possibility is to minimise the Kullback–Leibler divergence from the prior predictive distribution to the target
$\tilde{D}(\lambda\mid\boldsymbol{X})=\text{KL}(\text{T}(Y\mid\boldsymbol{X})\;\|\;\text{P}(Y\mid\lambda,\boldsymbol{X})).$ (19)
For discrete $Y$ the challenge remains, as when minimising the CvM and AD discrepancies, estimating $\text{P}(Y\mid\lambda,\boldsymbol{X})$; when $Y$ is continuous we instead require an estimate of $\text{p}(Y\mid\lambda,\boldsymbol{X})$; for mixed discrete-continuous cases, a suitable KL divergence definition is less obvious.
Suppose $\text{T}(Y\mid\boldsymbol{X})$ is multivariate Gaussian with mean $\mu_{1}$ and covariance $\boldsymbol{\Sigma}_{1}$ and, for a suitable range/value of $\lambda$, the prior predictive is well-approximated by another multivariate Gaussian $\hat{\text{P}}(Y)$ with mean $\hat{\mu}_{2}$ and covariance $\hat{\boldsymbol{\Sigma}}_{2}$. Given the assumption that we can draw samples from the prior predictive, this approximation is always possible. In this case, the KL divergence from $\hat{\text{P}}$ to T is
$\text{KL}(\text{T}(Y\mid\boldsymbol{X})\;\|\;\hat{\text{P}}(Y))=\frac{1}{2}\left[\log\frac{|\hat{\boldsymbol{\Sigma}}_{2}|}{|\boldsymbol{\Sigma}_{1}|}-d+\text{tr}\{\hat{\boldsymbol{\Sigma}}_{2}^{-1}\boldsymbol{\Sigma}_{1}\}+(\hat{\mu}_{2}-\mu_{1})^{T}\hat{\boldsymbol{\Sigma}}_{2}^{-1}(\hat{\mu}_{2}-\mu_{1})\right].$ (20)
For completeness, we also implement the “reverse” KL divergence $\text{KL}(\hat{\text{P}}(Y)\;\|\;\text{T}(Y\mid\boldsymbol{X}))$ (the direction denoted in Equation 20 is referred to as the “forward” KL divergence). We assess the differences between these KL-based discrepancies and the CvM discrepancy in human height growth example, where the Gaussian approximation to the prior predictive distribution is most appropriate.
S4.1 Ensuring computational equivalence when using the KL as a discrepancy
Our KL approximation in Equation 20 means we do not need to estimate an ECDF or perform numerical integration to compute the discrepancy. To ensure fair comparisons between this KL-based discrepancy and the Cramér-von Mises or Anderson-Darling discrepancies, the Gaussian approximation to the prior predictive distribution uses the same number of samples as the CvM and AD discrepancies use in their ECDF estimate of $\text{P}(Y\mid\lambda,\boldsymbol{X})$. For generality, we do not require the end-user to supply $\mu_{1}$ and $\boldsymbol{\Sigma}_{1}$. We instead estimate these parameters using samples from $\text{T}(Y\mid\boldsymbol{X})$, that would be employed in locating an appropriate importance sampling distribution. This approximation is available in our pbbo R package.
S5 Additional information for the cure fraction survival example
@@
S5.1 Hyperparameter support $\Lambda$
@@
S5.2 Pareto Frontiers for Cramér-von Mises discrepancy
@@
S5.3 Final objective values
@@
S6 Additional information for the $R^{2}$ example
S6.1 Hyperparameter support $\Lambda$ – faithfulness experiment
@@
S6.2 Full faithfulness results
@@
Figure 16: As in Figure 7 (main text) but for all values of $(s_{1},s_{2})$ denoted in the facet panels titles. The performance of the regularised horseshoe is superior to the Dirichlet-Laplace, both of which are vast improvements over the Gaussian.
S6.3 A comparison to an asymptotic result
@@
S6.3.1 Results and analytic comparison
@@
S7 Additional information for the Preece-Baines example
S7.1 Hyperparameter support $\Lambda$
@@
S7.2 Details for $\text{T}(Y)$ and $\text{T}(Y\mid X_{r})$
@@
S7.4 Choosing $\kappa^{*}$
Optimal values of $\kappa$ are selected for the multi-objective approaches by minimising the variance of the sum of both objectives. These values are displayed in Table 7, and are used for all multi-objective results in this section. The range values considered, $\mathcal{K}$, is specific to each target/discrepancy pair, as each objective is scale-free and thus universally applicable fixed ranges are not available.
Table 7: Selected optimal values $\kappa^{*}$ for each combination of target and discrepancy. Optimal values are selected as those that minimise the variance of the sum of both objectives.
Target Discrepancy $\kappa^{*}$ $\mathcal{K}$
Covariate-independent Cramér-von Mises 0.42 $[0.05,1.9]$
Covariate-independent KL forward 1.43 $[0.05,3.5]$
Covariate-independent KL reverse 0.28 $[0.05,6.0]$
Covariate-dependent Cramér-von Mises 0.24 $[0.05,0.5]$
Covariate-dependent KL forward 0.49 $[0.05,0.7]$
Covariate-dependent KL reverse 0.46 $[0.05,2.0]$
S7.5 Pareto frontiers for the covariate-independent target
The Pareto frontiers for the covariate-independent target for optimal $\kappa^{*}\in\mathcal{K}$, as defined in Table 7. is displayed in Figure 20.
@@
S7.6 Final predictive discrepancy values for the human growth example
Figure 21 displays the value of the discrepancy (CvM or KL, as appropriate) at the optima located by the multi-stage optimisation process. The optima are not comparable across targets, discrepancies (the KL-divergence is not a distance metric), and optimisation approaches, however for specific choices of these we can asses the variability across replicates, to eliminate incomplete-optimisation as a possible source of non-replicability/non-faithfulness. There is universally additional noise in the multiple objective approach, which is expected, and there is some slightly bi-modality in both KL-based discrepancies for the covariate-specific target. Overall, the multi-stage optimiser seems to consistently locate acceptable optima.
@@
Figure 21: Final predictive discrepancy $\log(D(\lambda^{*}))$, or $\log(D(\lambda^{*}\mid\boldsymbol{X}))$ for the covariate-specific target. The multiple objective optimisation approach uses the optimal values $\kappa^{*}$ listen in Table 7 of this supplement. Horizontal jitter has been applied for readability.
S7.7 Further assessing faithfulness
After asserting that optima are consistently found by our multi-stage optimisation process, we assess faithfulness by inspecting Figure 10 (in the main text) and Figure 22 in this supplement, which display the prior predictive and covariate-(in)dependent target distributions. All single-objective approaches are more faithful than their multi-objective counterparts, which is expected given we sacrifice some amount of faithfulness for variability in $\theta$ when using the multi-objective approach. Of the discrepancies, for the covariate-independent target (Figure 10, main text), the Cramér-von Mises discrepancy seems most faithful and replicable. For the covariate-specific target (Figure 22, this supplement), all discrepancies result in similarly faithful priors and prior predictive distributions. Both the single and multi-objective approaches struggle to match the prior predictive distribution at all ages, with consistently poorer faithfulness for $X_{1}=2$. Empirically, it does not seem possible to match all four margins of the supplied target prior predictive distributions simultaneously, given the mathematical structure of the model. Lastly, because $\text{t}(Y\mid X_{1}=2)$ is substantially narrower than the other targets, it is optimal, under the Cramér-Von Mises discrepancy and the forward KL, to select wider priors better matching the older age target distributions. The reverse KL is more concentrated than both the Cramér-von Mises and the forward KL, which is a known property of the KL in this direction (Minka, 2005). This concentration is also visible when inspecting the prior predictives for the conditional mean, $\text{p}(h(t;\theta)\mid\lambda^{*})$, displayed in Figure 11 of the main text.
Figure 22: The covariate-specific target densities $\text{t}(Y\mid\boldsymbol{X})$ (red) and prior predictive densities $\text{p}(Y\mid\lambda^{*},\boldsymbol{X})$ for each combination of discrepancy and single/multi-object approach, each of these with 30 replicates (blue lines). The columns depict the age-specific conditionals of this target.
S7.8 Further details of assessing prior replicability, uniqueness, and differences between KL and CvM discrepancies
We assess replicability and uniqueness in $\theta$ by inspecting $\text{p}(h_{0}\mid\lambda^{*})$ displayed in Figure 23. All target and discrepancy combinations seem to provide broadly replicable results when inspecting $h_{0}$, which agrees with the assessed optimisation convergence in Figure 21. However uniqueness remains an unsolved challenge, particularly for the covariate-specific target, where two distinct modes are visible across both single and multi-objective settings for all discrepancies. We highlight, for the covariate-specific target, the very similar marginal priors found using either the Cramér-von Mises, forward KL, or reverse KL discrepancy. This indicates, for at least this parameter and target, that the optimal prior is insensitive to the choice of discrepancy. We also observe wider priors for $h_{0}$ in the covariate-independent setting using the KL-based discrepancies, which further explains the implausibly flat (a priori) growth curves visible in Figure 11 of the main text.
Figure 23: The marginal prior $\text{p}(h_{0}\mid\lambda^{*})$ translated from both the covariate-independent and covariate-specific targets, for all discrepancies considered in this example. Multi-objective priors are chosen using the relevant value of $\kappa^{*}$ in Table 7.
S7.9 Full marginal prior and posterior comparison plots
@@
Figures 24 and 25 are extended versions of Figure 13 in the main text, and display the prior and posterior estimates for all the parameters in $\theta$. Note that results here are limited to the priors, and corresponding posteriors, obtained using the Cramér-Von Mises discrepancy. Consistency and uniqueness remain, evidently, challenging and as yet unobtainable.
@@
Figure 24: A comparison of the priors (blue) produced by our method using the covariate-independent marginal target (bottom two rows, Cramér-von Mises discrepancy only); and Hartmann et al. (2020) (second row), with no prior displayed for the flat prior scenario. The corresponding posteriors (red) for individual $n=26$ under each of these priors are displayed as dashed lines. Note that y-axes change within columns and are limited to values that clip some of the priors/posteriors for readability.